1) The molar solubility of silver sulfate in a 0.135 M ammonium sulfate solution is  M.

2) The molar solubility of lead phosphate in a 0.154 M lead acetate solution is  M.
Answer #1

According to HomeworkLib rules we have to answer only first one problem

1) molar solubility of silver sulfate :

Ksp = [Ag+]² [SO4]
Let molar solubility = x, so
Ksp = (2x)² (x)= 1.2 x 10^ -5

Given that 0.135 M ammonium sulfate

(NH4)2 SO4 = 2 NH4+   + SO4^2-

So SO4^2- = 0.135

Ksp = (2x)² (x + 0.44 M)
Make the same assumptions here as for part b.:
1.2 x 10^ -5 = (2x)² (0.135 M);

1.2 x 10^ -5/(0.135 M); = 4x²

4x²= 8.89*10^-5

x²=8.89*10^-5 /4

x²=2.23*10^-5 taking under root;

x= 4.71*10^-3 M